A. What is an equation?

B. Explain the difference between a linear and a quadratic equation in terms of degree and solution.

C. Give a concise definition of the following:
1. Rate
2. Ratio
3. Proportion
1
Expert's answer
2020-09-23T17:20:47-0400

Solution to A:

An equation is a statement of equality of two expressions formulated by applying various algebraic operations such as addition, subtraction, multiplication and division to a set of variables e.g

"3x^2 +4x-2=y"

Solution to B:

A linear equation is an equation where the highest degree of the unknown variable is one e.g "2x+4=6"

Whereas a quadratic equation is an equation where the highest degree of the unknown variable is two e.g "4x^2+3x-2=0"

A linear equation has only one solution whereas a quadratic equation has 2 solutions


Solution to C:

  1. Ratio is an expression of one measure relative to another e.g in a class with 40 boys and 37 girls, we can say the ratio of boys to girls in the class is 40:37. Ratio can be expressed either as a fraction "({40 \over 30} )" or in words (40 to 30) or using colon (40:30).
  2. Rate is a form of ratio that compares quantities of different types e.g speed is a rate that compares distance and time (km/hr)
  3. Proportion is a ratio that relates a part to a whole e.g the proportion of boys in the class above is "{40 \over 40+37} = 0.5195" or 51.95%